OpenStudy (anonymous):

What is the solution to the rational equation x over 5 plus 4 over 5 = 7 ?

OpenStudy (anonymous):

33? i have no idea. ive done it every way possible and just dont get it

OpenStudy (anonymous):

\[\frac{x}{5}+\frac{4}{5} = 7\]\[\frac{x}{5} = 7 - \frac{4}{5}\] Then try multiplying both sides by 5.

OpenStudy (anonymous):

what u mean both sides by five. like 4/5 (5) and x/5(5)

OpenStudy (anonymous):

\[5(\frac{x}{5})=5(7-\frac{4}{5})\]

OpenStudy (anonymous):

\[\frac{x}{5}+\frac{4}{5}=7\] multiply both sides by 5 \[x+4=35\] subtract 4 \[x=31\]

OpenStudy (anonymous):

31

OpenStudy (anonymous):

thanks u for clearing that up